# A Linearly Convergent Frank-Wolfe-type Method for Smooth Convex Minimization over the Spectrahedron

### v1

@@ Abstract
We consider the problem of minimizing a smooth and convex function over the $n$-dimensional spectrahedron — the set of real symmetric $n\times n$ positive semidefinite matrices with unit trace, which underlies numerous applications in statistics, machine learning and additional domains. Standard first-order methods often require high-rank matrix computations which are prohibitive when the dimension $n$ is large. The well-known Frank-Wolfe method on the other hand, only requires efficient rank-one matrix computations, however suffers from worst-case slow convergence, even under conditions that enable linear convergence rates for standard methods. In this work we present the first Frank-Wolfe-based algorithm that only applies efficient rank-one matrix computations and, assuming quadratic growth and strict complementarity conditions, is guaranteed, after a finite number of iterations, to converges linearly, in expectation, and independently of the ambient dimension.
@@ 1 Introduction
This paper is concerned with efficient first-order optimization methods (suitable for high dimensional settings) for minimizing a smooth (Lipschitz continuous gradient) and convex objective function $f:\mathbb{R}^{n\times n}\rightarrow\mathbb{R}$ over the $n$-dimensional spectrahedron — the set of all $n\times n$ real symmetric positive semidefinite matrices with unit trace $\mathcal{S}^{n}:=\{{\mathbf{X}}\in\mathbb{S}^{n}~{}|~{}{\mathbf{X}}\succeq 0,\textrm{Tr}({\mathbf{X}})=1\}$, where $\mathbb{S}^{n}$ denotes the space of real symmetric $n\times n$ matrices, and ${\mathbf{X}}\succeq 0$ denotes that ${\mathbf{X}}$ is positive semidefinite. This problem underlies numerous applications of interest in diverse fields of science and engineering such as statistics, machine learning, and discrete optimization, including for instance convex relaxations to low-rank matrix recovery problems [5, 26, 3, 4, 20, 28], covariance matrix estimation problems [29, 27, 6], convex relaxations to hard combinatorial problems [30, 15, 25], and many more . More specifically, we are interested in methods inspired by the well known Frank-Wolfe method (FW), aka the conditional gradient method [9, 22, 19]. These type of methods, while applicable more broadly to smooth convex minimization over arbitrary convex and compact subsets of an inner product space, are in particular interesting in case the feasible set is the spectrahedron, see for instance [17, 20, 11, 31, 10] 111some papers, e.g., [20, 10], consider the closely related problem of optimization over the matrix nuclear norm ball, see [20] for a discussion of the connection between these setups, on a number of accounts. First, while orthogonal projections onto the spectrahderon, that are applied in stanadard projection-based gradient methods, require in worst case a full eigen-decomposition of a symmetric $n\times n$ matrix, which in practical implementations runs in $O(n^{3})$ time, a linear optimization step over the spectrahedron as required by FW, amounts only to an efficient rank-one matrix computation: a single leading eigenvector computation of a symmetric $n\times n$ matrix, which often runs in nearly linear time in the size of the matrix using fast iterative eigenvector methods. Second, since FW only applies a rank-one update, in many cases of interest both the objective function value and gradient direction could be updated from one iteration to the next much more efficiently than when the update is of high rank. This is for instance the case when the objective $f$ includes terms such as $g(\mathcal{A}{\mathbf{X}})$ (e.g., in matrix sensing problems [31]) or $g(\mathcal{A}{\mathbf{X}}^{-1})$ (e.g., in covariance estimation problems [6]), where $\mathcal{A}:\mathbb{R}^{n\times n}\rightarrow\mathbb{R}^{n}$ is a linear map composed of rank-one matrices: $\mathcal{A}{\mathbf{X}}=({\mathbf{a}}_{1}^{\top}{\mathbf{X}}{\mathbf{b}}_{1},\dots,{\mathbf{a}}_{m}^{\top}{\mathbf{X}}{\mathbf{b}}_{m})^{\top}\in\mathbb{R}^{m}$ with $\{({\mathbf{a}}_{i},{\mathbf{b}}_{i})\}_{i=1}^{m}\subset\mathbb{R}^{n}\times\mathbb{R}^{n}$, and $g:\mathbb{R}^{m}\rightarrow\mathbb{R}$ is coordinate-wise separable. Another example, is when the objective function includes a $\log\det({\mathbf{X}})$ component. Third, another consequence of the low-rank updates of FW is that when initialized with a low-rank matrix, and provided that the number of iterations is substantially lower than the dimension $n$, the iterates produced by the method are low-rank and can be stored in factorized form to further reduce memory and runtime requirements. Finally, FW-based methods are often parameter-free and do not require tuning of any parameters (such as the smoothness or strong convexity parameters), and can often be implemented using simple and efficient line-search.
@@
The major drawback of Frank-Wolfe however is worst-case slow convergence rates. For minimizing a $\beta$-smooth convex objective over the spectrahedron, the worst-case number of iterations to guarantee an $\epsilon$-approximated solution in function value is $O(\beta/\epsilon)$ [17, 20, 19]. This does not improve even under standard curvature assumptions such quadratic growth (see for instance lower bound in [18]), which are well known to facilitate linear convergence rates (i.e., convergence rates that scale only with $\log(1/\epsilon)$) for standard projection-based methods [8, 24].
@@
In recent years, several Frank-Wolfe-inspired methods for optimization over the spectrahedron have been developed that, under conditions such as quadratic growth and/or strict complementarity, enjoy a linear convergence rate. [14] established that in case there exists a rank-one optimal solution which satisfies the strict complementarity condition, the standard FW with line-search, after a finite number of iteration, begins to converge linearly. This result however does not extend to the general case in which there is an optimal solution with rank greater than one. The works [1, 7] proposed methods, which we refer to as block-Frank-Wolfe methods, that converge linearly without restriction on the rank of optimal solutions, however these diverge significantly from the classical FW template: while standard FW only applies rank-one leading eigenvector computations, [1, 7] require on each iteration to compute the $r$ leading components in the singular value decomposition (SVD) of a $n\times n$ matrix, where the rank parameter $r$ is required to be at least as large as the highest rank of any optimal solution. As a result, these methods have several inherent and substantial limitations. First, they require knowledge of the rank of optimal solutions which is often unknown and overestimating this quantity results in needlessly expensive SVD computations. Second, these methods are only efficient when optimal solutions are of low-rank, since otherwise the partial SVD computations required are not significantly more efficient than those in projection-based methods. Similarly, in this case also the important FW feature of fast update times of objective function and gradient information is lost. Finally, while FW can often be implemented with simple line-search, the method in [1], aside of the smoothness constant $\beta$, requires the quadratic growth constant $\alpha$ in order to set the step-size, which is often very difficult to estimate. The method in [7] does not require these constants and instead performs a sophisticated multi-variate “step-size” optimization on each iteration, by minimizing the original objective function $f$ over an $r$-dimensional spectrahedron (where $r$ is the upper-bound on the rank of optimal solutions), for instance by running a nested first-order method. However, this is again may only be efficient for small values of $r$.
@@
We mention in passing that aside of linear convergence results, the work [11] presented a FW variant for the spectrahedron that also uses only rank-one matrix computations (leading eigenvector) and, under strong convexity of the objective function, is guaranteed to converge with a fast rate of the form $\min\{\frac{\textrm{rank}({\mathbf{X}}^{*})^{2/3}}{t^{4/3}},\frac{1}{\lambda_{\min}({\mathbf{X}}^{*})^{2}t^{2}}\}$, where $t$ is the iteration counter and $\lambda_{\min}({\mathbf{X}}^{*})$ denotes the smallest non-zero eigenvalue of the unique optimal solution ${\mathbf{X}}^{*}$. While this result does not limit the rank of the optimal solution and clearly improves upon the standard FW rate of $O(1/t)$, it is not a linear convergence rate.
@@
We answer this question in the negative, under the assumption that quadratic growth and strict complementarity conditions hold (see formal definitions in Section 1.2). We propose a novel Frank-Wolfe-based method that is guaranteed, after a finite number of iterations (“burn-in” phase), to converge linearly, in expectation (the expectation is due to the use of randomization in one of the steps of the algorithm). In particular, both the finite number of initial steps and the linear convergence rate are independent of the ambient dimension. Importantly, our method admits an implementation, which aside of three efficient leading eigenvector computations (the same type of rank-one computations applied in the standard FW method and can be executed in parallel), requires only $O(n^{2})$ runtime per iteration. In terms of parameter tuning, our method requires only knowledge of the smoothness constant of the objective function $\beta$ (which as opposed to the rank of optimal solutions or the quadratic growth constant, is often relatively easy to bound).
@@
This paper $1$ $\beta$ YES linear (in expectation)
@@
Before moving on, we pause to discuss the strict complementarity assumption in some more detail. While such a condition does not typically appear in linear convergence results for projection-based methods, it was in fact highly instrumental to obtaining linear convergence rates for FW-based methods in a broader context: It was introduced in the classical text [22] for optimization over strongly convex sets, in [16, 12] for obtaining dimension-independent linear rates for polytopes, and in [14, 7] for the spectrahedron setting. Moreover, [7] proved that in our spectrahedron setting with an objective function which admits the highly popular structure $f({\mathbf{X}})=g(\mathcal{A}{\mathbf{X}})$, where $\mathcal{A}$ is a linear map and $g$ is smooth and strongly convex, strict complementarity further implies that $f$ satisfies quadratic growth over the spectrahderon. [32] further provided numerical evidence for a closely related setting (though not completely identical to ours), that for an objective function with the above structure, linear convergence may be unattainable when strict complementarity does not hold.
@@ 1.1 Inspiration from the polytope setting
Aside from our matrix spectrahedron setting, the Frank-Wolfe method has gained significant interest in the past decades also for optimization over convex and compact polytopes. While in this setting also the worst-case convergence of the standard method (with line-search or without) scales with $1/\epsilon$, even under favorable conditions such as quadratic growth or strong convexity of the objective function, it was established in a series of works that simple modifications of the method, all of which rely in some way or the other on the concept of incorporating away steps into the algorithm, linear convergence can be established, and in particular that a linear rate independent of the ambient dimension could be established when strict complementarity also holds. In the celebrated work [16], which was the first to obtain a linear convergence result for polytopes (under strong convexity and strict complementarity), the authors proved that after a finite number of steps, all iterates of the method must lie inside the optimal face of the polytope, and there it acts as if performing unconstrained minimization and thus, once inside the optimal face, the strong convexity assumption implies linear convergence.
Our method and analysis are highly inspired by [16], but with some considerable differences. The main difficulty is that while polytopes have a finite number of faces which indeed allows to argue that under strict complementarity, from some iteration, all iterates will lie inside the optimal face, such an argument is not sensible for the spectraehedron which has infinitely many faces222a face of $\mathcal{S}^{n}$ is given by the set $\{{\mathbf{V}}{\mathbf{S}}{\mathbf{V}}^{\top}~{}|~{}{\mathbf{S}}\in\mathcal{S}^{r}\}$ for some integer $r\in\{1,\dots,n\}$ and matrix ${\mathbf{V}}\in\mathbb{R}^{n\times r}$ such that ${\mathbf{V}}^{\top}{\mathbf{V}}={\mathbf{I}}$. To bypass this difficulty, we apply our arguments w.r.t. to a time-changing face of the spectrahedron, one that is defined according to a principle subspace of the current gradient direction. Once close enough to an optimal solution (which takes a finite number of steps), we can distinguish between two main cases: if our current iterate is sufficiently aligned with this face, we can apply similar arguments to those in [16] (i.e., essentially unconstrained minimization) and argue that either a standard Frank-Wolfe step or a specialized away-step (designed specifically for our spectrahedron setting) will reduce the approximation error by a constant fraction. Otherwise (in case the current iterate is not sufficiently aligned with the current face), we establish that a specialized randomized pairwise step, i.e., a step that replaces a random rank-one component supported by the current iterate with a new rank-one component, reduces the error by a constant fraction, in expectation. Thus, an overall linear convergence rate (in expectation) is obtained.
@@ 1.2 Notation and Assumptions
[…]
We let $\mathcal{X}^{*}\subseteq\mathcal{S}^{n}$ denote the set of minimizers of the objective function $f$ over the spectrahedron $\mathcal{S}^{n}$, and we let $f^{*}$ denote the minimal value of $f$ over $\mathcal{S}^{n}$.
Recall we assume that $f$ is convex, and we also assume it is $\beta$-smooth over $\mathcal{S}^{n}$, i.e., there exists $\beta\geq 0$ such that for any ${\mathbf{X}},{\mathbf{Y}}\in\mathcal{S}^{n}$ we have that $\|{\nabla{}f({\mathbf{X}})-\nabla{}f({\mathbf{Y}})}\|_{F}\leq\beta\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{F}$. Recall this further implies the well known inequality (see for instance Lemma 5.7 in [2]):
@@
Our second assumption is strict complementarity, which as discussed in the Introduction, has played a central role in proving linear convergence rates for Frank-Wolfe-type methods in diverse settings. In case of the spectrahedron it takes the form of a positive eigengap in the gradient direction at optimal points (see also [7, 14, 13]).
Assumption 2 (strict complementarity).
@@
$\min_{{\mathbf{X}}^{*}\in\mathcal{X}^{*}}\left({\lambda_{n-r^{*}}(\nabla{}f({\mathbf{X}}^{*}))-\lambda_{n-r^{*}+1}(\nabla{}f({\mathbf{X}}^{*}))}\right)=\delta.$ (2)
@@
Note that Assumption 2 implies that there exists ${\mathbf{V}}^{*}\in\mathbb{R}^{n\times r^{*}}$ such that ${\mathbf{V}}^{*\top}{\mathbf{V}}^{*}={\mathbf{I}}$ and $\mathcal{X}^{*}\subseteq\{{\mathbf{V}}^{*}{\mathbf{S}}{\mathbf{V}}^{*\top}~{}|~{}{\mathbf{S}}\in\mathcal{S}^{r^{*}}\}$, that is, $\mathcal{X}^{*}$ lies on a $r^{*}$-dimensional face of $\mathcal{S}^{n}$.

### v2

# arXiv:2503.01441v2[math.OC] 02 Mar 2026

@@ Abstract
We consider the problem of minimizing a smooth and convex function over the $n$-dimensional spectrahedron — the set of real symmetric $n\times n$ positive semidefinite matrices with unit trace, which underlies numerous applications in statistics, machine learning and additional domains. Standard first-order methods often require high-rank matrix computations which are prohibitive when the dimension $n$ is large. The well-known Frank-Wolfe method on the other hand only requires efficient rank-one matrix computations, however, suffers from worst-case slow convergence, even under conditions that enable linear convergence rates for standard methods. In this work we present the first Frank-Wolfe-based algorithm that only applies efficient rank-one matrix computations and, assuming quadratic growth and strict complementarity conditions, is guaranteed, after a finite number of iterations, to converge linearly, in expectation, and independently of the ambient dimension.
@@ 1 Introduction
This paper is concerned with efficient first-order optimization methods (suitable for high dimensional settings) for minimizing a smooth (Lipschitz continuous gradient) and convex objective function $f:\mathbb{R}^{n\times n}\rightarrow\mathbb{R}$ over the $n$-dimensional spectrahedron — the set of all $n\times n$ real symmetric positive semidefinite matrices with unit trace $\mathcal{S}^{n}:=\{{\mathbf{X}}\in\mathbb{S}^{n}~|~{\mathbf{X}}\succeq 0,\textrm{Tr}({\mathbf{X}})=1\}$, where $\mathbb{S}^{n}$ denotes the space of real symmetric $n\times n$ matrices, and ${\mathbf{X}}\succeq 0$ denotes that $\mathbf{X}$} is positive semidefinite. This problem underlies numerous applications of interest in diverse fields of science and engineering such as statistics, machine learning, and discrete optimization, including for instance convex relaxations to low-rank matrix recovery problems [6, 27, 4, 5, 19, 29], covariance matrix estimation problems [30, 28, 7], convex relaxations to hard combinatorial problems [31, 16, 26], and many more. More specifically, we are interested in methods inspired by the well-known Frank-Wolfe method (FW), aka the conditional gradient method [10, 23, 21]. These types of methods, while applicable more broadly to smooth convex minimization over arbitrary convex and compact subsets of an inner product space, are in particular interesting in case the feasible set is the spectrahedron, see for instance [18, 19, 12, 32, 11] 111some papers, e.g., [19, 11], consider the closely related problem of optimization over the matrix nuclear norm ball, see [19] for a discussion of the connection between these setups, for a number of reasons. First, while orthogonal projections onto the spectrahedron, that are applied in standard projection-based gradient methods, require in worst case a full eigen-decomposition of a symmetric $n\times n$ matrix, which in practical implementations runs in $O(n^{3})$ time, a linear optimization step over the spectrahedron as required by FW, amounts only to an efficient rank-one matrix computation: a single leading eigenvector computation of a symmetric $n\times n$ matrix, which often runs in nearly linear time in the size of the matrix using fast iterative eigenvector methods. Second, since FW only applies a rank-one update, in many cases of interest both the objective function value and gradient direction could be updated from one iteration to the next much more efficiently than when the update is of high rank. This is for instance the case when the objective $f$ includes terms such as $g(\mathcal{A}{\mathbf{X}})$ (e.g., in matrix sensing problems [32]) or $g(\mathcal{A}{\mathbf{X}}^{-1})$ (e.g., in covariance estimation problems [7]), where $\mathcal{A}:\mathbb{R}^{n\times n}\rightarrow\mathbb{R}^{m}$ is a linear map composed of rank-one matrices: $\mathcal{A}{\mathbf{X}}=({\mathbf{a}}_{1}^{\top}{\mathbf{X}}{\mathbf{b}}_{1},\dots,{\mathbf{a}}_{m}^{\top}{\mathbf{X}}{\mathbf{b}}_{m})^{\top}\in\mathbb{R}^{m}$ with $\{({\mathbf{a}}_{i},{\mathbf{b}}_{i})\}_{i=1}^{m}\subset\mathbb{R}^{n}\times\mathbb{R}^{n}$, and $g:\mathbb{R}^{m}\rightarrow\mathbb{R}$ is coordinate-wise separable. Another example is when the objective function includes a $\log\det({\mathbf{X}})$ component. Third, another consequence of the low-rank updates of FW is that when initialized with a low-rank matrix, and provided that the number of iterations is substantially lower than the dimension $n$, the iterates produced by the method are low-rank and can be stored in factorized form to further reduce memory and runtime requirements. Finally, FW-based methods are often parameter-free and do not require tuning of any parameters (such as the smoothness or strong convexity parameters), and can often be implemented using simple and efficient line-search.
@@
The major drawback of Frank-Wolfe, however, is worst-case slow convergence rates. For minimizing a $\beta$-smooth convex objective over the spectrahedron, the worst-case number of iterations to guarantee an $\epsilon$-approximated solution in function value is $O(\beta/\epsilon)$ [18, 19, 21]. This does not improve even under standard curvature assumptions such as quadratic growth (see for instance lower bound in [20]), which are well-known to facilitate linear convergence rates (i.e., convergence rates that scale only with $\log(1/\epsilon)$) for standard projection-based methods [9, 25].
@@
In recent years, several Frank-Wolfe-inspired methods for optimization over the spectrahedron have been developed that, under conditions such as quadratic growth and/or strict complementarity, enjoy a linear convergence rate. [15] established that in case there exists a rank-one optimal solution which satisfies the strict complementarity condition, the standard FW with line-search, after a finite number of iterations, begins to converge linearly. This result, however, does not extend to the general case in which there is an optimal solution with rank greater than one. The works [1, 8] proposed methods, which we refer to as block-Frank-Wolfe methods, that converge linearly without restriction on the rank of optimal solutions, however these diverge significantly from the classical FW template: while standard FW only applies rank-one leading eigenvector computations, [1, 8] require on each iteration to compute the $r$ leading components in the singular value decomposition (SVD) of an $n\times n$ matrix, where the rank parameter $r$ is required to be at least as large as the highest rank of any optimal solution. As a result, these methods have several inherent and substantial limitations. First, they require knowledge of the rank of optimal solutions which is often unknown and overestimating this quantity results in needlessly expensive SVD computations. Second, these methods are only efficient when optimal solutions are of low-rank, since otherwise the partial SVD computations required are not significantly more efficient than those in projection-based methods. Similarly, in this case also the important FW feature of fast update times of objective function and gradient information is lost. Finally, while FW can often be implemented with simple line-search, the method in [1], aside from the smoothness constant $\beta$, requires the quadratic growth constant $\alpha$ in order to set the step-size, which is often very difficult to estimate. The method in [8] does not require these constants and instead performs a sophisticated multivariate “step-size” optimization on each iteration, by minimizing the original objective function $f$ over an $r$-dimensional spectrahedron (where $r$ is the upper-bound on the rank of optimal solutions), for instance by running a nested first-order method. However, this may again only be efficient for small values of $r$.
@@
We mention in passing that aside from linear convergence results, the work [12] presented a FW variant for the spectrahedron that also uses only rank-one matrix computations (leading eigenvector) and, under strong convexity of the objective function, is guaranteed to converge with a fast rate of the form $\min\{\frac{\textrm{rank}({\mathbf{X}}^{*})^{2/3}}{t^{4/3}},\frac{1}{\lambda_{\min}({\mathbf{X}}^{*})^{2}t^{2}}\}$, where $t$ is the iteration counter and $\lambda_{\min}({\mathbf{X}}^{*})$ denotes the smallest non-zero eigenvalue of the unique optimal solution $\mathbf{X}$$^{*}$. While this result does not limit the rank of the optimal solution and clearly improves upon the standard FW rate of $O(1/t)$, it is not a linear convergence rate.
@@
We answer this question in the negative, under the assumption that quadratic growth and strict complementarity conditions hold (see formal definitions in Section 1.2). We propose a novel Frank-Wolfe-based method that is guaranteed, after a finite number of iterations (“burn-in” phase), to converge linearly, in expectation (the expectation is due to the use of randomization in one of the steps of the algorithm). In particular, both the finite number of initial steps and the linear convergence rate are independent of the ambient dimension. Importantly, our method admits an implementation, which aside from three efficient leading eigenvector computations (the same type of rank-one computations applied in the standard FW method and can be executed in parallel), requires only $O(n^{2})$ runtime per iteration. In terms of parameter tuning, our method requires only knowledge of the smoothness constant of the objective function $\beta$ (which as opposed to the rank of optimal solutions or the quadratic growth constant, is often relatively easy to bound).
@@
This paper 11 $\beta$ YES linear (in expectation)
@@
Before moving on, we pause to discuss the strict complementarity assumption in some more detail. While such a condition does not typically appear in linear convergence results for projection-based methods, it was in fact highly instrumental to obtaining linear convergence rates for FW-based methods in a broader context: It was introduced in the classical text [23] for optimization over strongly convex sets, in [17, 13] for obtaining dimension-independent linear rates for polytopes, and in [15, 8] for the spectrahedron setting. Moreover, [8] proved that in our spectrahedron setting with an objective function which admits the highly popular structure $f({\mathbf{X}})=g(\mathcal{A}{\mathbf{X}})$, where $\mathcal{A}$ is a linear map and $g$ is smooth and strongly convex, strict complementarity further implies that $f$ satisfies quadratic growth over the spectrahedron. [33] further provided numerical evidence for a closely related setting (though not completely identical to ours), that for an objective function with the above structure, linear convergence may be unattainable when strict complementarity does not hold.
@@ 1.1 Inspiration from the polytope setting
Aside from our matrix spectrahedron setting, the Frank-Wolfe method has gained significant interest in the past decades also for optimization over convex and compact polytopes. While in this setting also the worst-case convergence of the standard method (with line-search or without) scales with $1/\epsilon$, even under favorable conditions such as quadratic growth or strong convexity of the objective function, it was established in a series of works that simple modifications of the method, all of which rely in some way or another on the concept of incorporating away steps into the algorithm, linear convergence can be established, and in particular that a linear rate independent of the ambient dimension could be established when strict complementarity also holds. In the celebrated work [17], which was the first to obtain a linear convergence result for polytopes (under strong convexity and strict complementarity), the authors proved that after a finite number of steps, all iterates of the method must lie inside the optimal face of the polytope, and there it acts as if performing unconstrained minimization and thus, once inside the optimal face, the strong convexity assumption implies linear convergence.
Our method and analysis are highly inspired by [17], but with some considerable differences. The main difficulty is that while polytopes have a finite number of faces which indeed allows to argue that under strict complementarity, from some iteration, all iterates will lie inside the optimal face, such an argument is not sensible for the spectrahedron which has infinitely many faces [^2] A face of $\mathcal{S}^{n}$ is given by the set $\{{\mathbf{V}}{\mathbf{S}}{\mathbf{V}}^{\top}~|~{\mathbf{S}}\in\mathcal{S}^{r}\}$ for some integer $r\in\{1,\dots,n\}$ and matrix ${\mathbf{V}}\in\mathbb{R}^{n\times r}$ such that ${\mathbf{V}}^{\top}{\mathbf{V}}={\mathbf{I}}$. To bypass this difficulty, we apply our arguments w.r.t. a time-changing face of the spectrahedron, one that is defined according to a principal subspace of the current gradient direction. Once close enough to an optimal solution (which takes a finite number of steps), we can distinguish between two main cases: if our current iterate is sufficiently aligned with this face, we can apply similar arguments to those in [17] (i.e., essentially unconstrained minimization) and argue that either a standard Frank-Wolfe step or a specialized away-step (designed specifically for our spectrahedron setting) will reduce the approximation error by a constant fraction. Otherwise (in case the current iterate is not sufficiently aligned with the current face), we establish that a specialized randomized pairwise step, i.e., a step that replaces a random rank-one component supported by the current iterate with a new rank-one component, reduces the error by a constant fraction, in expectation. Thus, an overall linear convergence rate (in expectation) is obtained.
@@ 1.2 Notation and Assumptions
We denote matrices by uppercase boldface letters, e.g., $\mathbf{X}$}, column vectors by lowercase boldface letters, e.g., $\mathbf{v}$}, and scalars are denoted by lightface letters, e.g., $r,\alpha,C$. For matrices we let $\|{\cdot}\|$ denote the spectral norm (largest singular value) and for column vectors we let it denote the Euclidean norm. We let $\|{\cdot}\|_{F}$ denote the Frobenius (Euclidean) norm for matrices. For a matrix ${\mathbf{A}}\in\mathbb{S}^{n}$ we let $\lambda_{1}({\mathbf{A}}),\dots,\lambda_{n}({\mathbf{A}})$ denote its (real) eigenvalues in non-ascending order. We denote by $\textrm{Im}({\mathbf{A}})$ and $\textrm{Ker}({\mathbf{A}})$ its image and kernel, respectively, and by $\mathbf{A}$$^{\dagger}$ its Moore-Penrose pseudoinverse. We let $\mathbf{I}$} denote the identity matrix whose dimension will be clear from context. We let $\mathbb{S}^{n}_{+}$ denote the positive semidefinite cone in $\mathbb{S}^{n}$. We let $\langle{{\mathbf{A}},{\mathbf{B}}}\rangle=\textrm{Tr}({\mathbf{A}}^{\top}{\mathbf{B}})$ denote the standard matrix inner-product.
We let $\mathcal{X}^{*}\subseteq\mathcal{S}^{n}$ denote the set of minimizers of the objective function $f$ over the spectrahedron $\mathcal{S}^{n}$, and we let $f^{*}$ and $\nabla{}f^{*}$ denote the minimal value of $f$ over $\mathcal{S}^{n}$ and the gradient direction at any minimizer (note that since $f$ is differentiable, the gradient direction is constant over $\mathcal{X}^{*}$).
Recall we assume that $f$ is convex, and we also assume it is $\beta$-smooth over $\mathcal{S}^{n}$, i.e., there exists $\beta\geq 0$ such that for any ${\mathbf{X}},{\mathbf{Y}}\in\mathcal{S}^{n}$ we have $\|{\nabla{}f({\mathbf{X}})-\nabla{}f({\mathbf{Y}})}\|_{F}\leq\beta\|{{\mathbf{X}}-{\mathbf{Y}}}\|_{F}$. Recall this further implies the well-known inequality (see for instance Lemma 5.7 in [2]):
@@
Our second assumption is that all optimal solutions satisfy the strict complementarity condition, which as discussed in the Introduction, has played a central role in proving linear convergence rates for Frank-Wolfe-type methods in diverse settings. In the case of the spectrahedron this takes the form of a positive eigengap in the gradient direction at optimal solutions (see also [8, 15, 14]) and that all optimal solutions have the same rank.
Assumption 2 (strict complementarity of optimal solutions).
@@
$\lambda_{n-r^{*}}(\nabla{}f^{*})-\lambda_{n-r^{*}+1}(\nabla{}f^{*})=\delta.$ (2)
@@
Note Assumption 2 implies that there exists ${\mathbf{V}}^{*}\in\mathbb{R}^{n\times r^{*}}$ such that ${\mathbf{V}}^{*\top}{\mathbf{V}}^{*}={\mathbf{I}}$ and $\mathcal{X}^{*}\subset\{{\mathbf{V}}^{*}{\mathbf{S}}{\mathbf{V}}^{*\top}~|~{\mathbf{S}}\in\mathcal{S}^{r^{*}},{\mathbf{S}}\succ 0\}$, that is, $\mathcal{X}^{*}$ lies in the interior of a $r^{*}$-dimensional face of $\mathcal{S}^{n}$. This follows from the first-order optimality condition, see for instance Lemma 5.2 in [14].
One may wonder if in order to obtain our results it is indeed required, in case the optimal solution is not unique, that all of them satisfy strict complementarity (which mandates that all solutions have the same rank). In some previous analyses of Frank-Wolfe methods with strict complementarity conditions [17, 13, 15] it was assumed that the optimal solution is unique (e.g., by assuming $f$ is strongly convex). Our analysis, which does not assume uniqueness, indeed requires that all solutions satisfy this condition. It remains open whether this is mandatory.